A company is mixing a blend of two different coffees. The first kind (x) costs $8 a pound, and the second (y) costs $5 per pound. How much of each should they use if they want 60 pounds worth $6.25 per pound?
Answer by jorel1380(3719)

You have two different coffees. If the first is x, and the second is y, then y=60-x. So:
8x+5(60-x)=6.25(60)
3x+300=375
3x=75
x=25
y=35
The company wants 25 lbs. of the first coffee for every 35 lbs. of the second. ☺☺☺☺